Determine the number that will complete the square to solve the equation after the constant term has been written on the right side. Do not solve the equation.
3x^2-20x-1=0
The number is ____.

Answer by rothauserc(4718)

3x^2 -20x -1 = 0
:
add 1 to both sides of =
:
3x^2 -20x = 1
:
divide the coefficient of the x term by 2 and square it, then add it to both sides of =
:
3x^2 -20x +100 = 1 +100 = 101
:
*********************************************************************
the number that will complete the square to solve the equation is 101
*********************************************************************
: